OpenStudy (anonymous):

find the limit as x -->2 7/x-2

OpenStudy (anonymous):

would it be 0/0 or undefined

OpenStudy (anonymous):

equation is 7/x-2

OpenStudy (anonymous):

you can also find the limit by derivating the numerator and denominator the derivatives of 7 is 0 and the derivatives of x-2 is 1 therefore the limit would be 0/1 = 0

OpenStudy (anonymous):

IF you're asking about \[\lim_{x \rightarrow 2}{7 \over x-2}\], then you can notice that when substituting \(x=2\), we get a \(0\) at the denominator. Therefore, the limit is undefined and goes to \(\infty\) as x approaches \(2^+\) and \(-\infty\) as x approaches \(2^-\).

OpenStudy (anonymous):

AnwarA has the correct approach. So in this case, L'Hopital's rule can't be used to evaluate the limit, as the third post states.